A body is dropped from a certain height h and second body is thrown downwards with velocity of 5m /s simultaneously . What will be the differences in heights of the two bodies after 3sec

Grade:11

1 Answers

Hasan Naqvi
8 years ago
Distance covered by first ball = 0 + ½ * 10 * 32= 45
 
Distance covered by second ball=5 x 3 + ½ x 10 x 33 = 15 + 45 = 60
 
Difference in height = 15m